I'm having a really hard time figuring this problem out please help
What is the area in feet of circular tablecloth that has a diameter of 42 inches?

Answer by RAY100(1637)

Remember that Area = pi * rad^2
.
we are given diameter = 42, or radius = 21 ,,,,(42/2)
.
Area = pi * r^2 = pi * 21^2 = pi * 441 = 1385.44 inches ^2
.
.
an obscure Area formula is A = pi d^2 /4 = pi 42^2 /4 = 1385.44 also
.